1.05 grams of a nonvolatile nonionizing solute is added to 625 mL water. The osmotic pressure of the resulting solution is 0.269 atm at 28.2 °C. What is the molar mass of the unknown?
Answer #1

Let n be the no.of moles of solute.

Hence, Molarity of the solution, M = (n)/(0.625)

Osmotic pressure, P = 0.269 atm

Temperature, T = 273 + 28.2 K = 301.2 K

Substituting in the equation of osmotic pressure:

Therefore, Molar mass of the solute will be: